### v2

@@ Abstract
In this paper, we determine the leading order term of $\vartheta(G(n,k,L))$ of any generalized Johnson graph with $k$ and $L$ fixed and $n\rightarrow\infty$. As an application of this theorem, we give an explicit construction of a graph $G$ on $n$ vertices with a large gap between the Lovász number and the Shannon capacity $c(G)$. Specifically, we prove that for any $\epsilon>0$, for sufficiently large $n$ there is a generalized Johnson graph $G$ on $n$ vertices which has ratio $\vartheta(G)/c(G)=\Omega(n^{1-\epsilon})$, which improves on all known constructions. The graph $G$ a fortiori also has ratio $\vartheta(G)/\alpha(G)=\Omega(n^{1-\epsilon})$, which greatly improves on the best known explicit construction.
@@ 1 Introduction
The correspondence between $L$-systems and independent sets of generalized Johnson graphs $G(n,k,L)$ implies that if $\mathcal{F}\subset\binom{[n]}{k}$ is an $L$-system, then $|\mathcal{F}|\leq\alpha(G(n,k,L))$. In this paper, we study the Lovász number of generalized Johnson graphs $G(n,k,L)$. The Lovász number $\vartheta(G)$ is a semidefinite programming approximation for the independence number $\alpha(G)$ of a graph $G$. The Lovász number is useful to compute the Shannon capacity $c(G)$ [^1] The Shannon capacity is more typically denoted $\Theta(G)$. We use the notation $c(G)$ to avoid clashes with the asymptotic notation $\Theta$.. The Shannon capacity [Sha56] of a graph $G$ is defined as $c(G)=\lim_{n\rightarrow\infty}(\alpha(G_{n}))^{1/n}$, where $G_{n}=\alpha(G^{\boxtimes n})$ (recall the strong product $\boxtimes$ of two graphs $G$ and $H$ is the graph $G\boxtimes H$ with vertex set $V(G)\times V(H)$, and $(u_{1},v_{1})\sim(u_{2},v_{2})$ if and only if $u_{1}=u_{2}$ and $v_{1}\sim v_{2}$ or $u_{1}\sim u_{2}$ and $v_{1}=v_{2}$ or $u_{1}\sim u_{2}$ and $v_{1}\sim v_{2}$). The Shannon capacity is notoriously hard to compute, as there is no known algorithm which computes the Shannon capacity in a finite amount of time; in fact, there isn’t even a known algorithm which can approximate the Shannon capacity to within a constant factor. Lovász [Lov79] proved that $\alpha(G)\leq c(G)\leq\vartheta(G)$, so if it can be shown that $\vartheta(G)=\alpha(G)$ (as Lovász [Lov79, Theorem 13] showed for the Kneser graph), then the Shannon capacity of the graph $G$ is also determined. Indeed, Schrijver [Sch81] (for $n\rightarrow\infty$) and Wilson [Wil84] (for the optimal range $n\geq(t+1)(k-t+1)$) proved the $t$-intersecting Erdős-Ko-Rado theorem by showing $\vartheta(G(n,k,[t,k-1]))=\binom{n-t}{k-t}$, so it is natural to wonder how good an upper bound for $\alpha(G(n,k,L))$ can be obtained from $\vartheta(G(n,k,L))$. For example, does $\vartheta(G(n,k,L))$ give the correct order of magnitude for $\alpha(G(n,k,L))$?
The first main result of this paper is the determination of the leading order term of $\vartheta(G(n,k,L))$ for fixed $k$ and $L$ with $n\rightarrow\infty$. Let $L=\{\ell_{1},\ell_{2},\ldots,\ell_{s}\}$ with $\ell_{1}<\ell_{2}<\ldots<\ell_{s}$. A set of integers $\{\ell_{m},\ldots,\ell_{m+p}\}\subset L$ is a run of consecutive integers in $L$ if $\ell_{m+i}=\ell_{m}+i$ for every $0\leq i\leq p$. The set $\{\ell_{m},\ldots,\ell_{m+p}\}$ is a full run of consecutive integers in $L$ if additionally the following two conditions hold: (1) either $m=1$ or $\ell_{m-1}<\ell_{m}-1$; (2) either $m+p=s$ or $\ell_{m+p+1}>\ell_{m+p}+1$. The set $L$ may be divided uniquely into full runs of consecutive integers. For example, if $L=\{1,3,4,7,8,9,11\}$, then $L=\{1\}\cup\{3,4\}\cup\{7,8,9\}\cup\{11\}$ contains $4$ runs of consecutive integers. The length of a run is the cardinality of the set associated with the run.
@@
Let $G=G(n,k,L)$ be a generalized Johnson graph for some integers $n>k>0$ and a set $L=\{\ell_{1},\ell_{2},\ell_{3},\ldots,\ell_{s}\}\subset[0,k-1]$ with $\ell_{1}<\ell_{2}<\ldots<\ell_{s}$. Suppose $L$ contains $b$ full runs of consecutive integers. Then, as $n\rightarrow\infty$, there is a constant $c$ depending on $k$ and $L$ but not $n$ such that
@@
and a constant $C$ depending on $k,L$ and $L^{C}:=[0,k-1]\setminus L$ but not $n$ such that
@@
so the Lovász number bound does not do worse than Theorem 1.3 for at least one of $L$ and $L^{C}$. Schrijver’s question is equivalent to asking if $\vartheta(G(n,k,L))=\prod_{\ell\in L}\frac{n-\ell}{k-\ell}$ as $n\rightarrow\infty$. Theorem 1.2 shows that this is not the case for most choices of $k$ and $L$.
@@
$\vartheta((G(n,k,L))=1+\frac{\binom{k}{k-\ell}\binom{n-k}{k-\ell}}{\sum_{j=0}^{k-\ell}(-1)^{j+1}\binom{\ell+1}{j}\binom{k-\ell-1}{k-\ell-j}\binom{n-k-\ell-1}{k-\ell-j}}=\frac{\binom{k}{k-\ell}}{(k-\ell)(\ell+1)}n+O(1).$
@@
$\vartheta(G(n,k,L^{\prime}))=\frac{\binom{n}{k}}{\vartheta((G(n,k,L))}=\frac{(\ell+1)(k-\ell)}{k!\binom{k}{k-\ell}}n^{k-1}+O(n^{k-2}).$
@@
On the other hand, we can use Theorem 1.2 to give explicit examples of graphs with large gaps between the Shannon capacity and the Lovász number (and a fortiori large gaps between the independence number and the Lovász number). The question of how large the ratio $\vartheta(G)/\alpha(G)$ can be for a graph on $n$ vertices is related to the question of how good an approximation $\vartheta(G)$ is for the independence number and the Shannon capacity. Lovász noted (see [Knu94]) that for random graphs $\vartheta(G)/\alpha(G)=\Theta(n^{\frac{1}{2}}/\log n)$, while Feige [Fei97] gave a randomized construction showing that there are graphs on $n$ vertices with $\alpha(G)=n^{o(1)}$ and $\vartheta(G)=n^{1-o(1)}$. However, there does not seem to be much known about explicit constructions which give fairly large values of $\vartheta(G)/\alpha(G)$ or $\vartheta(G)/c(G)$. Alon and Kahale [AK98] showed that for any $\epsilon>0$, there is a Frankl-Rödl graph on $n$ vertices with $\vartheta(G)\geq(\frac{1}{2}-\epsilon)n$ and $\alpha(G)=O(n^{\delta})$, where $\delta=\delta(\epsilon)<1$. Peeters [Pee96, Remark 2.1] showed that the symplectic graphs $\text{Sp}(2m,2)$ have $c(\text{Sp}(2m,2))=\log_{2}(n+1)+1$, while $\vartheta(\text{Sp}(2m,2))=\sqrt{n+1}+1$, implying a ratio of $\vartheta/c=\Theta(n^{1/2}/\log n)$ (the number of vertices for these graphs is $n=2^{2m}-1)$.
Our second main result is that for any $\epsilon>0$, there is an explicit construction of a graph $G$ on $n$ vertices with ratio $\vartheta(G)/c(G)=\Omega(n^{1-\epsilon})$. This also implies large ratios for $\vartheta(G)/\alpha(G)$ and $\chi(G)/\vartheta(G^{c})$.
@@
For any $\epsilon>0$ and $n$ sufficiently large, there is an explicit construction of a graph $G$ on $n$ vertices with:
@@ 2.2 Analyzing the linear program for the Lovász number of generalized Johnson graphs
$\frac{1}{\binom{k}{k-\ell}\binom{n-k}{k-\ell}}\sum_{j=0}^{k-\ell}(-1)^{j}\binom{u}{j}\binom{k-u}{k-\ell-j}\binom{n-k-u}{k-\ell-j}=O(1),$
@@
$\frac{1}{\binom{k}{k-\ell}\binom{n-k}{k-\ell}}\sum_{j=0}^{k-\ell}(-1)^{j}\binom{u}{j}\binom{k-u}{k-\ell-j}\binom{n-k-u}{k-\ell-j}=O\left(\frac{1}{n^{u-\ell}}\right),$
@@
for some constant $c$ depending on $k$ and $L$ but not $n$. This is sufficient to prove Theorem 1.2, as the same argument with $L^{C}=[0,k-1]\setminus\{\ell_{1},\ell_{2},\ldots,\ell_{s}\}:=\{\ell_{1}^{\prime},\ell_{2}^{\prime},\ldots,\ell_{k-s}^{\prime}\}$ implies that
$\vartheta(G(n,k,L^{C}))\geq\prod_{i=1}^{b^{\prime}}m_{i}!\cdot\frac{\binom{k}{k-\ell_{1}^{\prime}}\binom{k-\ell_{1}^{\prime}-1}{k-\ell_{2}^{\prime}}\cdots\binom{k-\ell_{k-s-1}^{\prime}-1}{k-\ell_{k-s}^{\prime}}}{\prod_{i=1}^{k-s}(\ell_{i}^{\prime}+1)(k-\ell_{i}^{\prime})}n^{k-s}-n^{k-s-1},$
@@
which together complete the proof of the claim. Indeed, consider two consecutive entries $\ell_{i}^{\prime}$ and $\ell_{i+1}^{\prime}$ in $L^{\prime}$. If $\ell_{i+1}^{\prime}=\ell_{i}^{\prime}+1$, then $(\ell_{i+1}^{\prime}-\ell_{i}^{\prime}-1)!=0!=1$. Otherwise, the integers $[\ell_{i}^{\prime}+1,\ell_{i+1}^{\prime}-1]$ correspond to a full run of consecutive integers in $L$ of length $(\ell_{i+1}^{\prime}-1)-(\ell_{i}^{\prime}+1)+1=\ell_{i+1}^{\prime}-\ell_{i}^{\prime}-1$, and each full run of consecutive integers in $L$ can be obtained in this way. The same argument with full runs in $L^{\prime}$ and consecutive entries in $L$ gives the other equality. ∎
@@
Let $P_{s}$ be the $s\times{s}$ matrix which has $(i,j)$-entry $P_{s_{(i,j)}}=P_{k-\ell_{j}}^{\ell_{i}+1}$, so that
$P_{s}=\begin{pmatrix}P_{k-\ell_{1}}^{\ell_{1}+1}&P_{k-\ell_{2}}^{\ell_{1}+1}&\ldots&P_{k-\ell_{s}}^{\ell_{1}+1}\\ \vdots&\vdots&&\vdots\\ P_{k-\ell_{1}}^{\ell_{s}+1}&P_{k-\ell_{2}}^{\ell_{s}+1}&\ldots&P_{k-\ell_{s}}^{\ell_{s}+1}\end{pmatrix}.$
@@
$P_{s}{\bf a}\geq\begin{pmatrix}-1\\ -1\\ \vdots\\ -1\end{pmatrix}.$
@@
${\bf v}=P_{s}^{-1}\begin{pmatrix}0\\ 0\\ \vdots\\ 0\\ -1\end{pmatrix}.$
It is not a priori clear that such a ${\bf v}$ exists, but we will show subsequently in Lemma 2.8 that $P_{s}$ is invertible as $n\rightarrow\infty$. We claim that as $n\rightarrow\infty$, the following is a feasible solution to this linear program:
@@
Claim 2.7 has two important consequences. First, we use Claim 2.7 to show that $P_{s}$ is invertible as $n\rightarrow\infty$. In fact, we determine the leading order term of $\text{det}(P_{s})$.
@@
$\text{det}(P_{s})=(-1)^{s}Cn^{sk-(\ell_{1}+\ell_{2}+\cdots+\ell_{s})-s}+O(n^{sk-(\ell_{1}+\cdots+\ell_{s})-s-1}),$
@@
In particular, $P_{s}$ is invertible as $n\rightarrow\infty$ and the solution (4) exists.
@@
$P_{s_{(i,j)}}=P_{k-\ell_{j}}^{\ell_{i}+1}=\begin{cases}\frac{1}{(k-\ell_{j})!}\binom{k-\ell_{i}-1}{k-\ell_{j}}n^{k-\ell_{j}}-O(n^{k-\ell_{j}-1}),&i<j\\ (-1)^{\ell_{i}+1-\ell_{j}}\binom{\ell_{i}+1}{\ell_{j}}\frac{1}{(k-\ell_{i}-1)!}n^{k-\ell_{i}-1}+(-1)^{\ell_{i}-\ell_{j}}O(n^{k-\ell_{i}-2}),&i\geq j.\end{cases}$ (5)
@@
is the set of inversions of $\sigma$. By (5), in row $i$ the highest order of magnitude term is of the order $n^{k-\ell_{i}-1}$, so the highest order terms of $\text{det}(P_{s})$ are of the order $n^{\sum_{i=1}^{s}(k-\ell_{i}-1)}=n^{sk-(\ell_{1}+\cdots+\ell_{s})-s}$. Let $\Sigma$ be the set of all permutations $\sigma\in S_{s}$ which have the property that
$\prod_{i=1}^{s}P_{s_{(i,\sigma(i))}}=\Theta(n^{sk-(\ell_{1}+\ell_{2}+\cdots+\ell_{s})-s}).$
These are the highest order terms in the Leibniz formula for $\text{det}(P_{s})$. We show the following result, which will prove Lemma 2.8.
$\sum_{\sigma\in\Sigma}\text{sgn}(\sigma)\prod_{i=1}^{s}P_{s_{(i,\sigma(i))}}=(-1)^{s}\frac{\prod_{i=1}^{s}(\ell_{i}+1)}{\prod_{i=1}^{b}m_{i}!\prod_{i=1}^{s}(k-\ell_{i}-1)!}n^{sk-(\sum_{i=1}^{s}\ell_{i})-s}+O(n^{sk-(\sum_{i=1}^{s}\ell_{i})-s-1}).$ (6)
@@
Let $\sigma\in\Sigma$. Then, for any $11\leq i\leq s-$ 1, $\sigma(i)\leq i+1$. Furthermore, if $i$ is the last integer in its full run of consecutive integers in $L$, then $\sigma(i)\leq i$.
@@
Let $\sigma\in S_{s}$ be a permutation with $\sigma(i)\geq i+2$ for some $21\leq i\leq s-$ 2. Then, $P_{s_{(i,\sigma(i))}}=O(n^{k-\ell_{i+2}})=o(n^{k-\ell_{i}-1})$, so the term in the Leibniz product formula corresponding to $\sigma$ will be of an order smaller than $n^{sk-(\ell_{1}+\ell_{2}+\cdots+\ell_{s})-s}$.
Similarly, if $\sigma$ is a permutation with $\sigma(i)\geq i+1$, where $i$ is an integer which is last in its full run of consecutive integers, then $P_{s_{(i,\sigma(i))}}=O(n^{k-\ell_{i+1}})=o(n^{k-\ell_{i}-1})$, since $\ell_{i+1}\geq\ell_{i}+2$, so again the term in the Leibniz product formula corresponding to $\sigma$ will be of smaller order. ∎
Claim 2.9 implies that every $\sigma\in\Sigma$ can only permute elements within each full run; that is, there are no $i$ and $j$ in different full runs of consecutive integers in $L$ with $\sigma(i)=j$. We will therefore consider runs of consecutive integers in $L$. Let $L_{m}=\{\ell_{m+1},\ell_{m+2},\ldots,\ell_{m+p}\}$ be a (not necessarily full) run of $p$ consecutive integers in $L$ with $\ell_{m+j+1}=\ell_{m+j}+1$ for $11\leq j\leq p-$ 1. We prove a result on the terms of the matrix $P_{s}$ coming from $L_{m}$ which will ultimately allow us to prove (6).

### v3

@@ Abstract
In this paper, we determine the leading order term of $\vartheta(G(n,k,L))$ of any generalized Johnson graph with $k$ and $L$ fixed and $n\rightarrow\infty$. As an application of this theorem, we give an explicit construction of a graph $G$ on $n$ vertices with a large gap between the Lovász number and the Shannon capacity $c(G)$. Specifically, we prove that for any $\epsilon>0$, for infinitely many $n$ there is a generalized Johnson graph $G$ on $n$ vertices which has ratio $\vartheta(G)/c(G)=\Omega(n^{1-\epsilon})$, which improves on all known constructions. The graph $G$ a fortiori also has ratio $\vartheta(G)/\alpha(G)=\Omega(n^{1-\epsilon})$, which greatly improves on the best known explicit construction.
@@ 1 Introduction
The correspondence between $L$-systems and independent sets of generalized Johnson graphs $G(n,k,L)$ implies that if $\mathcal{F}\subset\binom{[n]}{k}$ is an $L$-system, then $|\mathcal{F}|\leq\alpha(G(n,k,L))$. In this paper, we study the Lovász number of generalized Johnson graphs $G(n,k,L)$. The Lovász number $\vartheta(G)$ is a semidefinite programming approximation for the independence number $\alpha(G)$ of a graph $G$. The Lovász number is useful to compute the Shannon capacity $c(G)$ [^1] The Shannon capacity is more typically denoted $\Theta(G)$. We use the notation $c(G)$ to avoid clashes with the asymptotic notation $\Theta$.. The Shannon capacity [Sha56] of a graph $G$ is defined as $c(G)=\lim_{n\rightarrow\infty}(\alpha(G_{n}))^{1/n}$, where $G_{n}=G^{\boxtimes n}$ (recall the strong product $\boxtimes$ of two graphs $G$ and $H$ is the graph $G\boxtimes H$ with vertex set $V(G)\times V(H)$, and $(u_{1},v_{1})\sim(u_{2},v_{2})$ if and only if $u_{1}=u_{2}$ and $v_{1}\sim v_{2}$ or $u_{1}\sim u_{2}$ and $v_{1}=v_{2}$ or $u_{1}\sim u_{2}$ and $v_{1}\sim v_{2}$). The Shannon capacity is notoriously hard to compute, as there is no known algorithm which computes the Shannon capacity in a finite amount of time; in fact, there isn’t even a known algorithm which can approximate the Shannon capacity to within a constant factor. Lovász [Lov79] proved that $\alpha(G)\leq c(G)\leq\vartheta(G)$, so if it can be shown that $\vartheta(G)=\alpha(G)$ (as Lovász [Lov79, Theorem 13] showed for the Kneser graph), then the Shannon capacity of the graph $G$ is also determined. Indeed, Schrijver [Sch81] (for $n\rightarrow\infty$) and Wilson [Wil84] (for the optimal range $n\geq(t+1)(k-t+1)$) proved the $t$-intersecting Erdős-Ko-Rado theorem by showing $\vartheta(G(n,k,[t,k-1]))=\binom{n-t}{k-t}$, so it is natural to wonder how good an upper bound for $\alpha(G(n,k,L))$ can be obtained from $\vartheta(G(n,k,L))$. For example, does $\vartheta(G(n,k,L))$ give the correct order of magnitude for $\alpha(G(n,k,L))$?
The first main result of this paper is the determination of the leading order term of $\vartheta(G(n,k,L))$ for fixed $k$ and $L$ with $n\rightarrow\infty$. Let $L=\{\ell_{1},\ell_{2},\ldots,\ell_{s}\}$ with $\ell_{1}<\ell_{2}<\ldots<\ell_{s}$. A set of integers $\{\ell_{m},\ldots,\ell_{m+p}\}\subset L$ is a run of consecutive integers in $L$ if $\ell_{m+i}=\ell_{m}+i$ for every $0\leq i\leq p$. The set $\{\ell_{m},\ldots,\ell_{m+p}\}$ is a full run of consecutive integers in $L$ if additionally the following two conditions hold: (1) either $m=1$ or $\ell_{m-1}<\ell_{m}-1$; (2) either $m+p=s$ or $\ell_{m+p+1}>\ell_{m+p}+1$. The set $L$ may be divided uniquely into full runs of consecutive integers. For example, if $L=\{1,3,4,7,8,9,11\}$, then $L=\{1\}\cup\{3,4\}\cup\{7,8,9\}\cup\{11\}$ contains $4$ full runs of consecutive integers. The length of a run is the cardinality of the set associated with the run.
@@
Let $G=G(n,k,L)$ be a generalized Johnson graph for some integers $n>k>0$ and a set $L=\{\ell_{1},\ell_{2},\ell_{3},\ldots,\ell_{s}\}\subset[0,k-1]$ with $\ell_{1}<\ell_{2}<\ldots<\ell_{s}$. Suppose $L$ contains $b$ full runs of consecutive integers. Then, there is a constant $c$ depending on $k$ and $L$ but not on $n$ such that
@@
and a constant $C$ depending on $k$ and $L$, but not on $n$ such that
@@
so the Lovász number bound does not do worse than Theorem 1.3 for at least one of $L$ and $L^{C}$. Schrijver’s question is equivalent to asking if $\vartheta(G(n,k,L))=\prod_{\ell\in L}\frac{n-\ell}{k-\ell}$ as $n\rightarrow\infty$. Theorem 1.2 gives the leading order term of $\vartheta(G(n,k,L))$ for all $k$ and $L$ and allows one to easily construct many examples of $k$ and $L$ for which $\vartheta(G(n,k,L))\neq\prod_{\ell\in L}\frac{n-\ell}{k-\ell}$.
@@
$\vartheta(G(n,k,L))=1+\frac{\binom{k}{k-\ell}\binom{n-k}{k-\ell}}{\sum_{j=0}^{k-\ell}(-1)^{j+1}\binom{\ell+1}{j}\binom{k-\ell-1}{k-\ell-j}\binom{n-k-\ell-1}{k-\ell-j}}=\frac{\binom{k}{k-\ell}}{(k-\ell)(\ell+1)}n+O(1).$
@@
$\vartheta(G(n,k,L^{\prime}))=\frac{\binom{n}{k}}{\vartheta(G(n,k,L))}=\frac{(\ell+1)(k-\ell)}{k!\binom{k}{k-\ell}}n^{k-1}+O(n^{k-2}).$
@@
On the other hand, we can use Theorem 1.2 to give explicit examples of graphs with large gaps between the Shannon capacity and the Lovász number (and a fortiori large gaps between the independence number and the Lovász number). The question of how large the ratio $\vartheta(G)/\alpha(G)$ can be for a graph on $n$ vertices is related to the question of how good an approximation $\vartheta(G)$ is for the independence number and the Shannon capacity. Lovász noted (see [Knu94]) that for random graphs $\vartheta(G)/\alpha(G)=\Theta(n^{\frac{1}{2}}/\log n)$, while Feige [Fei97] gave a randomized construction showing that there is an infinite family of graphs with $\alpha(G)=n^{o(1)}$ and $\vartheta(G)=n^{1-o(1)}$. However, there does not seem to be much known about explicit constructions which give fairly large values of $\vartheta(G)/\alpha(G)$ or $\vartheta(G)/c(G)$. Alon and Kahale [AK98] showed that for any $\epsilon>0$, there is a Frankl-Rödl graph on $n$ vertices with $\vartheta(G)\geq(\frac{1}{2}-\epsilon)n$ and $\alpha(G)=O(n^{\delta})$, where $\delta=\delta(\epsilon)<1$. Peeters [Pee96, Remark 2.1] showed that the symplectic graphs $\text{Sp}(2m,2)$ have $c(\text{Sp}(2m,2))=\log_{2}(n+1)+1$, while $\vartheta(\text{Sp}(2m,2))=\sqrt{n+1}+1$, implying a ratio of $\vartheta/c=\Theta(n^{1/2}/\log n)$ (the number of vertices for these graphs is $n=2^{2m}-1)$.
Our second main result is that for any $\epsilon>0$, for infinitely many $n$ there is an explicit construction of a graph $G$ on $n$ vertices with ratio $\vartheta(G)/c(G)=\Omega(n^{1-\epsilon})$. This also implies large ratios for $\vartheta(G)/\alpha(G)$ and $\chi(G)/\vartheta(G^{c})$.
@@
For any $\epsilon>0$ and infinitely many $n$, there is an explicit construction of a graph $G$ on $n$ vertices with:
@@ 2.2 Analyzing the linear program for the Lovász number of generalized Johnson graphs
$\frac{1}{\binom{k}{k-\ell}\binom{n-k}{k-\ell}}\sum_{j=0}^{k-\ell}(-1)^{j}\binom{u}{j}\binom{k-u}{k-\ell-j}\binom{n-k-u}{k-\ell-j}=\frac{\binom{k-u}{k-\ell}}{\binom{k}{k-\ell}}-O\left(\frac{1}{n}\right)=O(1),$
@@
$\frac{1}{\binom{k}{k-\ell}\binom{n-k}{k-\ell}}\sum_{j=0}^{k-\ell}(-1)^{j}\binom{u}{j}\binom{k-u}{k-\ell-j}\binom{n-k-u}{k-\ell-j}=(-1)^{u-\ell_{1}}O\left(\frac{1}{n^{u-\ell}}\right),$
@@
for some constant $c$ depending on $k$ and $L$ but not on $n$. This is sufficient to prove Theorem 1.2, as the same argument with $L^{C}=[0,k-1]\setminus\{\ell_{1},\ell_{2},\ldots,\ell_{s}\}:=\{\ell_{1}^{\prime},\ell_{2}^{\prime},\ldots,\ell_{k-s}^{\prime}\}$ implies that
$\vartheta(G(n,k,L^{C}))\geq\prod_{i=1}^{b^{\prime}}m_{i}!\cdot\frac{\binom{k}{k-\ell_{1}^{\prime}}\binom{k-\ell_{1}^{\prime}-1}{k-\ell_{2}^{\prime}}\cdots\binom{k-\ell_{k-s-1}^{\prime}-1}{k-\ell_{k-s}^{\prime}}}{\prod_{i=1}^{k-s}(\ell_{i}^{\prime}+1)(k-\ell_{i}^{\prime})}n^{k-s}-c^{\prime}n^{k-s-1},$
@@
which together complete the proof of the claim. Indeed, consider two consecutive entries $\ell_{i}^{\prime}$ and $\ell_{i+1}^{\prime}$ in $L^{\prime}$ (and for the purpose of the following argument, set $\ell_{0}^{\prime}=-1$ and $\ell_{k-s+1}^{\prime}=k$). If $\ell_{i+1}^{\prime}=\ell_{i}^{\prime}+1$, then $(\ell_{i+1}^{\prime}-\ell_{i}^{\prime}-1)!=0!=1$. Otherwise, the integers $[\ell_{i}^{\prime}+1,\ell_{i+1}^{\prime}-1]$ correspond to a full run of consecutive integers in $L$ of length $(\ell_{i+1}^{\prime}-1)-(\ell_{i}^{\prime}+1)+1=\ell_{i+1}^{\prime}-\ell_{i}^{\prime}-1$, and each full run of consecutive integers in $L$ can be obtained in this way. The same argument with full runs in $L^{\prime}$ and consecutive entries in $L$ gives the other equality. ∎
@@
Let $P$ be the $s\times{s}$ matrix which has $(i,j)$-entry $P_{(i,j)}=P_{k-\ell_{j}}^{\ell_{i}+1}$, so that
$P=\begin{pmatrix}P_{k-\ell_{1}}^{\ell_{1}+1}&P_{k-\ell_{2}}^{\ell_{1}+1}&\ldots&P_{k-\ell_{s}}^{\ell_{1}+1}\\ \vdots&\vdots&&\vdots\\ P_{k-\ell_{1}}^{\ell_{s}+1}&P_{k-\ell_{2}}^{\ell_{s}+1}&\ldots&P_{k-\ell_{s}}^{\ell_{s}+1}\end{pmatrix}.$
@@
$P{\bf a}\geq\begin{pmatrix}-1\\ -1\\ \vdots\\ -1\end{pmatrix}.$
@@
${\bf v}=P^{-1}\begin{pmatrix}0\\ 0\\ \vdots\\ 0\\ -1\end{pmatrix}.$
It is not a priori clear that such a ${\bf v}$ exists, but we will show subsequently in Lemma 2.8 that $P$ is invertible as $n\rightarrow\infty$. We claim that as $n\rightarrow\infty$, the following is a feasible solution to this linear program:
@@
Claim 2.7 has two important consequences. First, we use Claim 2.7 to show that $P$ is invertible as $n\rightarrow\infty$. In fact, we determine the leading order term of $\text{det}(P)$.
@@
$\text{det}(P)=(-1)^{s}Cn^{sk-(\ell_{1}+\ell_{2}+\cdots+\ell_{s})-s}+O(n^{sk-(\ell_{1}+\cdots+\ell_{s})-s-1}),$
@@
In particular, $P$ is invertible as $n\rightarrow\infty$ and the solution (4) exists.
@@
$P_{(i,j)}=P_{k-\ell_{j}}^{\ell_{i}+1}=\begin{cases}\frac{1}{(k-\ell_{j})!}\binom{k-\ell_{i}-1}{k-\ell_{j}}n^{k-\ell_{j}}-O(n^{k-\ell_{j}-1}),&i<j\\ (-1)^{\ell_{i}+1-\ell_{j}}\binom{\ell_{i}+1}{\ell_{j}}\frac{1}{(k-\ell_{i}-1)!}n^{k-\ell_{i}-1}+(-1)^{\ell_{i}-\ell_{j}}O(n^{k-\ell_{i}-2}),&i\geq j.\end{cases}$ (5)
@@
is the set of inversions of $\sigma$. By (5), in row $i$ the highest order of magnitude term is of the order $n^{k-\ell_{i}-1}$, so the highest order terms of $\text{det}(P)$ are of the order $n^{\sum_{i=1}^{s}(k-\ell_{i}-1)}=n^{sk-(\ell_{1}+\cdots+\ell_{s})-s}$. Let $\Sigma$ be the set of all permutations $\sigma\in S_{s}$ which have the property that
$\prod_{i=1}^{s}P_{(i,\sigma(i))}=\Theta(n^{sk-(\ell_{1}+\ell_{2}+\cdots+\ell_{s})-s}).$
These are the highest order terms in the Leibniz formula for $\text{det}(P)$. We show the following result, which will prove Lemma 2.8.
$\sum_{\sigma\in\Sigma}\text{sgn}(\sigma)\prod_{i=1}^{s}P_{(i,\sigma(i))}=(-1)^{s}\frac{\prod_{i=1}^{s}(\ell_{i}+1)}{\prod_{i=1}^{b}m_{i}!\prod_{i=1}^{s}(k-\ell_{i}-1)!}n^{sk-(\sum_{i=1}^{s}\ell_{i})-s}+O(n^{sk-(\sum_{i=1}^{s}\ell_{i})-s-1}).$ (6)
@@
Let $\sigma\in\Sigma$. Then, for any $11\leq i\leq s-$ 1, $\sigma(i)\leq i+1$. Furthermore, if $\ell_{i}$ is the last integer in its full run of consecutive integers in $L$, then $\sigma(i)\leq i$.
@@
Let $\sigma\in S_{s}$ be a permutation with $\sigma(i)\geq i+2$ for some $21\leq i\leq s-$ 2. Then, $P_{(i,\sigma(i))}=O(n^{k-\ell_{i+2}})=o(n^{k-\ell_{i}-1})$, so the term in the Leibniz product formula corresponding to $\sigma$ will be of an order smaller than $n^{sk-(\ell_{1}+\ell_{2}+\cdots+\ell_{s})-s}$.
Similarly, if $\sigma$ is a permutation with $\sigma(i)\geq i+1$, where $\ell_{i}$ is an integer which is last in its full run of consecutive integers, then $P_{(i,\sigma(i))}=O(n^{k-\ell_{i+1}})=o(n^{k-\ell_{i}-1})$, since $\ell_{i+1}\geq\ell_{i}+2$, so again the term in the Leibniz product formula corresponding to $\sigma$ will be of smaller order. ∎
Claim 2.9 implies that every $\sigma\in\Sigma$ can only permute elements within each full run; that is, there are no $\ell_{i}$ and $\ell_{j}$ in different full runs of consecutive integers in $L$ with $\sigma(i)=j$. We will therefore consider runs of consecutive integers in $L$. Let $L_{m}=\{\ell_{m+1},\ell_{m+2},\ldots,\ell_{m+p}\}$ be a (not necessarily full) run of $p$ consecutive integers in $L$ with $\ell_{m+j+1}=\ell_{m+j}+1$ for $11\leq j\leq p-$ 1. We prove a result on the terms of the matrix $P$ coming from $L_{m}$ which will ultimately allow us to prove (6).